OpenStudy (anonymous):

Shirley invested a total of $12,000 in two certificates of deposit. One pays 6% interest, and the other pays 7% interest. Her total interest at the end of one year is $790. How much is invested in the CD that pays 6% interest?

OpenStudy (anonymous):

if you put say \(x\) as the amount invested at 6% then the amount invested at 7% must be \(12 00-x\) then \(.06\times x+.07(12000-x)=790\) solve for \(x\)

OpenStudy (anonymous):

probably easier to get rid of the annoying decimals by multiplying both sides by 100 and solve \[6x+7(12000-x)=79000\]